OpenStudy (anonymous):

Which of the following is the equation of the line that passes through (2, 6) and (-8, -14)? a. y = -2x + 2 b. y = 2x – 2 c. y = 2x + 2 d. y = -2x -2

OpenStudy (anonymous):

option a

OpenStudy (anonymous):

to write the equation - find the slope: m=(6-(-14))/(2-(-8))=20/10=2 now build your equation: y-6=2(x-2) y=2x+2 option C)